# Quantum Electrodynamics of graphene Landau levels in a deep-sub-wavelength hyperbolic phonon polariton cavity

### v1

@@ Abstract
The confinement of electromagnetic radiation within extremely small volumes offers an effective means to significantly enhance light-matter interactions, to the extent that zero-point quantum vacuum fluctuations can influence and control the properties of materials. Here, we develop a theoretical framework for the quantum electrodynamics of graphene Landau levels embedded in a deep-sub-wavelength hyperbolic cavity, where light is confined into ultrasmall mode volumes. By studying the spectrum, we discuss the emergence of polaritons, and disentangle the contributions of resonant quantum vacuum effects from those of purely electrostatic interactions. Finally, we study the hybridization between magnetoplasmons and the cavity’s electromagnetic modes.
@@ I Introduction
Human beings naturally use optical light to perceive the surrounding world through sight. Leveraging technological advancements, other light frequencies can be used to probe the properties of matter. For example, X-rays reveal the atomic structure of crystals Compton , and Raman spectroscopy Raman measures vibrational modes of molecules. After the seminal work of Purcell Purcell1946 , it is now well established that light can be tailored by means of resonant structures - or cavities - to enhance its interaction with matter. Indeed, the emergence of cavity quantum electrodynamics (C-QED) Raimond01 ; Haroche13 , where mirrors confine the electromagnetic field, has enabled the measurement of quantum properties of a single atom coupled to a single photon.
@@
Light-matter hybridization manifests as an avoided crossing between energy levels, known as the vacuum Rabi splitting $\Omega$Sanchez-Mondragon1983 ; agarwal1984 ; agarwal1985 ; Brune1996 ; Raimond01 , which in solid-state systems can be probed in transmission and reflection experimentsKhitrova2006 ; Todorov_PhysRevLett_2010 ; Todorov_PhysRevLett_2009 ; Scalari_Science_2012 ; Bitton . This phenomenon can be understood within both classical and quantum settingsDDB_arxiv_2023 but, in the context of C-QED, the quantum nature of the electric field $\hat{E}$ unveils a general correspondence between the vacuum Rabi splitting and quantum vacuum fluctuations of the electric field, $\Omega\sim\sqrt{\langle\hat{E}^{2}\rangle}$ dalibard1982 ; ciuti2005 ; DDB_arxiv_2023 . A promising strategy to enhance vacuum fluctuations (and consequently the coupling) is to confine the electromagnetic field to a small volume $V$chang2006 ; Kockum19 ; tame2013 , as the amplitude of fluctuations scales as $1/V$. While Fabry-Pérot cavities are often usedRaimond01 ; Haroche13 , the minimal achievable volume in these objects is limited by the so-called diffraction limit, $V\geq\lambda_{\rm d}^{3}$, where $\lambda_{\rm d}\equiv\lambda_{0}/2$ and $\lambda_{0}$ is the wavelength in free space. This limit imposes a significant obstacle to the compression and impact of quantum vacuum fluctuations.
Conversely, sub-wavelength cavities, made from an optically active electromagnetic environment, differ from optical cavities as they can confine light in a volume $V\ll\lambda_{\rm d}^{3}$, well below the diffraction limit chang2006 ; koppens2011 ; tame2013 ; bozhevolnyi2006 ; oulton2009 ; baumberg2019 ; tame2013 . Paradigmatic examples of sub-wavelength cavities are plasmonic systems and metamaterials Maier2007 . Plasmonic nanocavities exploit the excitation of surface plasmons at metal-dielectric interfaces to confine electromagnetic fields to volumes significantly smaller than the diffraction limit Schuller2010 ; Ciraci2012 ; Chikkaraddy2016 .
@@
While these systems can achieve extremely high compression factorsAlcaraz2018 ; Barra-Burillo2021 , they suffer from a low quality factor ($Q\approx 10$) stemming from the highly dissipative nature of metallic systems. Among these, split-ring resonators operating in the terahertz regimeMaissen2014 have been extensively used in solid-state cavities experiments, due to their integrability with high-quality two-dimensional electron gases. More generally, polaritonic systems serve as promising platforms for solid-state C-QED, as they can host deep sub-wavelength optical modes and offer great tunability dai2014 . While this approach is experimentally promising, most current theories are developed for Fabry-Pérot cavities, where excitations are described by a transverse vector potential $\hat{\bm{A}}$. These theories are applied to sub-wavelength cavities by using a phenomenological substitution to enhance the Rabi frequency with a compression factor $A=V/\lambda_{\rm d}^{3}$. The compression factor quantifies how much a sub-wavelength cavity surpasses the diffraction limit Keller2017 ; arwas2023 , reaching values of $A\approx 10^{-10}$. Given a Rabi frequency $\Omega$ derived from optical cavity theory, which scales as $\sim 1/\sqrt{V}$, the same quantity for a sub-wavelength cavity is obtained by substituting $\Omega\to\Omega/\sqrt{A}$, thus enhancing light-matter interactionsgao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 by a factor up to $10^{5}$. This phenomenological procedure underlies striking predictions in sub-wavelength cavities, such as cavity-mediated superconductivity schlawin2019 , breakdown of topological integer quantum Hall effect arwas2023 , and cavity control of topological properties wang2019 ; nguyen2023 . However, it has been shown andolina2024 that sub-wavelength cavities, characterized by a strong compression $A\ll 1$, cannot be described by a transverse vector potential $\hat{\bm{A}}$ but, rather, by a scalar potential $\hat{\phi}$, reflecting their electrostatic characterandolina2024 ; pantazopoulos2024 . This underscores the need for a comprehensive theory of sub-wavelength cavity QED based on the coupling with a scalar potential $\hat{\phi}$. Consequently, it is important to revisit the conclusions of Refs. gao2020, ; schlawin2019, ; wang2019, ; nguyen2023, ; zezyulin2022, ; lenk2022, ; arwas2023, ; pavosevic2022, in light of this new framework. Developing this theory is the main goal of the present article.
The excitations of a sub-wavelength cavity are not transverse photons but longitudinal hybrid light-matter excitations such as plasmonsGPlasmonic and polaritonsPCavities . Sub-wavelength cavities have recently been shown riolo2024 to alter the quasiparticle properties of Fermi liquids realized in weakly correlated 2D electron systems.
@@
Among sub-wavelength platforms, hyperbolic phonon-polariton (HPP) cavities dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore based on hexagonal boron nitride (hBN) offer an unparalleled balance of high compression factor ($A\approx 10^{-6}$) and impressive quality factors ($Q\approx 400$) HerzigSheinfux . Due to the strong anisotropy in its optical phonon spectrumcaldwell2014 , hBN exhibits an upper ($\sim 168-200$ meV) and a lower ($\sim 94-103$ meV) Reststrhalen band, where the in-plane and out-of-plane permittivity components are opposite in sign, leading to a hyperbolic polaritonic dispersion. The resulting ray-like nature of hBN mid-infrared phonon polaritons allows for extreme wavelength compression of light, enabling electromagnetic waves to be confined to deeply sub-wavelength scalesli2015 . On the other hand, the bulk nature of hBN phonon polaritons protects them from Ohmic loss mechanisms typical of metal-dielectric interfaces, enhancing the lifetime and making HPP cavities highly effective for mid-IR photonic applicationscaldwell2014 .
@@
In this paper, we present a fully non-local quantum theory to describe the cavity quantum electrodynamics of 2D electron systems coupled to sub-wavelength cavities, aiming to derive the vacuum Rabi frequencies from a fully microscopic theoretical framework without relying on any phenomenological assumptions. Specifically, we investigate electrons in graphene subject to a strong perpendicular magnetic field ${\bm{B}}$ and coupled to the sub-wavelength modes of HPP cavity. This ${\bm{B}}$ field induces circular electron motion (cyclotron motion), leading to the emergence of quantized energy levels known as Landau levels (LLs)Giuliani_and_Vignale ; PrangeGirvin1987 . The excitations of HPP cavities made from hBN are typically around $0.1~{}{\rm eV}$ HerzigSheinfux , hence LL transitions can be tuned to this energy range by applying a magnetic field of a few Tesla. Therefore, if the coupling is higher than the losses, LLs transitions hybridize with resonant cavity modes, forming Landau-phonon-polaritons (LPhPs) hagenmuller2010 ; hagenmuller2012 ; Chirolli2012 ; Pellegrino2014 ; ashida2023 .
@@
To describe LPhPs, we develop a comprehensive microscopic model of quantum vacuum fluctuations in deep sub-wavelength HPP cavities. With our microscopic theory, we can predict the optimal cavity size to maximize light-matter coupling, measured by the vacuum Rabi frequencies, without resorting to any phenomenological compression factor. Importantly, the deep sub-wavelength regime allows cavity and matter excitations to hybridize at very high wavevectors, necessitating a fully non-local theoretical framework. Indeed, the maximal vacuum Rabi frequencies are obtained for finite momenta. Scattering-type near-field optical microscopy (s-SNOM) basov2016 , capable of probing collective modes at such high wavevectors, provides a promising platform to measure our predictions. The fully microscopic theory developed here can have profound implications for the control of materials in sub-wavelength cavities. As a major example, it will be interesting in the future to use our theory to study quantum transport in these 2D electron systems embedded in HPP cavities. Several experiments showed that sub-wavelength split-ring cavities affect the integer appugliese2022 ; enkner2024 and fractional enkner2024b quantum Hall, generating significant theoretical interest ciuti2021 ; arwas2023 ; bacciconi2023 ; rokaj2023 ; rokaj2022 . In principle, our framework can be applied to study how HPP cavities modify magnetotransport paravicini2019 and topology in general arwas2023 ; nguyen2023 ; wang2019 ; Zeno ; Dmytruk2022 ; Dmytruk2024 ; Dag2024 . In our case, as the collective magnetic excitations are strongly coupled to the cavity, a fully many-body treatment of quantum transport would be needed, rather than a single-particle approach. Since in sub-wavelength cavities the light-matter coupling occurs via the density operator, rather than via the current, these cavities circumvent the so-called “no-go theorems”andolina_prb_2019 ; andolina_epjp_2022 ; rzazewski_prl_1975 ; nataf_naturecommun_2010 and thus can, in principle, realize a superradiant phase transition. However, since the excitations of the cavity are encoded in the scalar potential $\hat{\phi}$, the nature of the phase transition would be ferroelectricdebernardis2018 ; rouse2023 , and would not lead to photon condensation.
@@
This paper is organized as follows. In Section II, we summarize the main concepts and key results of this work, which are presented in greater detail in the subsequent sections. In Section III, we develop a quantum theory of HPP cavities. Section IV provides an overview of the physics of LLs in graphene. In Section V, we delve into the details of light-matter interactions and analyze the resulting polaritonic spectrum. Additional technical details regarding the electrostatic properties of hexagonal boron nitride are provided in Appendix A, while the density-density response function of Landau levels in graphene is reviewed in Appendix B.
@@ II Summary of main results
In this work, we present our primary findings structured as follows: First, in Sect. II.1, we solve the retarded Poisson equation within a sub-wavelength cavity to determine the scalar Green’s function. This solution enables us to quantize the scalar potential $\hat{\phi}$, expressing it in terms of the quantum modes of the cavity. This provides a theoretical framework for describing electromagnetic and quantum vacuum fluctuations within a subwavelength cavity. Thus, in Sect. II.2 we review the physics of LLs emerging in graphene subjected to a strong perpendicular magnetic field. Crucially, we highlight that, in sub-wavelength cavities, electrons couple through the density rather than the current as in traditional Fabry-Pérot cavities. Finally, in Sect. II.3 we develop the C-QED of a 2D material embedded in a sub-wavelength cavity. By formulating and diagonalizing the Hopfield Hamiltonian, we study polaritonic states that arise from the hybridization of LL transitions with cavity modes. This also allows us to directly calculate, without invoking any phenomenological procedure, the vacuum Rabi frequencies of the system, showing the system achieves ultra-strong and super-strong coupling,
While our theory is specifically applied to graphene embedded in an hBN cavity, the same theory can be readily applied to other 2D materials within sub-wavelength cavities.
@@ II.1 Theory of the Green’s function of HPP cavities
Generally, electromagnetic cavities are fully characterized by the solutions of Maxwell’s equationsJackson . In sub-wavelength cavities the most important contribution to the description of the electromagnetic properties is provided by the solution of the retarded Poisson equation in the presence of an external chargeandolina2024 ; riolo2024 :
@@
The cavity geometry that we consider is shown in Fig. 1. We further assume translational invariance along the $\hat{\bm{x}}$ and $\hat{\bm{y}}$ directions that denote 2D positions in this plane with the vector $\bm{r}_{\parallel}=x\hat{\bm{x}}+y\hat{\bm{y}}$. In the direction perpendicular to this plane, i.e. the $\hat{\bm{z}}$ direction and for $|z|\leq L_{z}/2$, we assume that space is filled by a hBN layer of thickness $L_{z}$ dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore , characterized by a uniaxial permittivity tensortomadin2015 ; riolo2024 $\bm{\epsilon}(\bm{r},\omega)=\text{diag}(\epsilon_{x}(\bm{r},\omega),\epsilon_{x}(\bm{r},\omega),\epsilon_{z}(\bm{r},\omega))$. This dielectric tensor is initially assumed to be real, with its imaginary part introduced later through the substitution $\omega\to\omega+\mathrm{i}\eta$. While this approach is not generally required, it is particularly advantageous in this context, as it allows us to quantize the field first and incorporate losses subsequently.
Further details on this tensor are reported in Appendix A. For $|z|>L_{z}/2$, we assume the presence of two perfect conductors, acting as mirrors. Consequently, the boundary conditions for the potential inside hBN ( $|z|\leq L_{z}/2$) are given by Jackson : $\phi(\bm{r}_{\parallel},z=\pm L_{z}/2)=0$.
@@
where $\bm{q}_{\parallel}$ is the wavevector in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane, ${q}_{\parallel}$ is its modulus, and $q_{z,n_{z}}\equiv(2n_{z}+1)\pi/L_{z}$. Eq. (LABEL:eqn:green_function_hBN_expansion_Main) represents a key theoretical result, as it shows that the scalar Green’s function can be expanded into distinct modes, which can then be quantized systematically. Notice that the Green’s function depends only on the modulus ${q}_{\parallel}$ as the system is invariant under rotations in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane. Fig. 2(a) displays the spectrum $-{\rm Im}~{}g^{(0)}({q}_{\parallel},0,0,\omega)$ of the HPP cavity, which quantifies the absorption of the cavity at the position $z=z^{\prime}=0$, where we will place the 2D material. The scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ exhibits poles when the condition ${\epsilon}_{z}(\omega)q_{z,n_{z}}^{2}+{\epsilon}_{x}(\omega)q_{\parallel}^{2}=0$ is satisfied. This occurs only in the lower ($\omega_{z}^{\rm T}<\omega<\omega_{z}^{\rm L}$) and upper ($\omega_{x}^{\rm T}<\omega<\omega_{x}^{\rm L}$) Reststrahlen bands. The longitudinal and transverse phonon frequencies, $\omega_{i}^{\rm L}$ and $\omega_{i}^{\rm T}$, are marked by black dotted lines in Fig. 2(a). We denote the poles of the bare scalar Green’s function with the symbol $\omega_{{q}_{\parallel},n_{z},s}$, where $s=+1$ ($s=-1$) refers to the upper (lower) Reststrahlen band. There are important differences and similarities between the scalar Green’s function in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) for HPP cavities and the full dyadic Green’s function for Fabry-Pérot cavities andolina2024 ; Charlie_Thesis . Both functions share key features, such as poles corresponding to mode frequencies and a prefactor of $1/L_{z}$, which encodes the compression mechanism. However, in both cases, the dependence on $L_{z}$ is non-trivial, as the mode frequencies $\omega_{{q}_{\parallel},n_{z},s}$ themselves depend on the cavity length $L_{z}$. In Fabry-P’erot cavities, the poles, which are associated with the energies of cavity photons, appear only in the vectorial part of the dyadic Green’s function andolina2024 , while the scalar part only represents the Coulomb potential modified by the mirrors’image charges. Conversely, in HPP cavities, poles corresponding to the phonon-polariton energies are present in both the scalar and vectorial components of the dyadic Green’s function. Nevertheless, due to the sub-wavelength nature of HPP cavity modes ($\lambda_{\rm p}=2\pi/q_{\parallel}\ll\lambda_{0}$) the scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ is the dominant contribution to the system’s electromagnetic properties andolina2024 .

### v2

# Quantum Electrodynamics of graphene Landau levels in a deep sub-wavelength hyperbolic phonon polariton cavity

@@ Abstract
The confinement of electromagnetic radiation within extremely small volumes offers an effective means to significantly enhance light-matter interactions, to the extent that zero-point quantum vacuum fluctuations can influence and control the properties of materials. Here, we develop a theoretical framework for the quantum electrodynamics of graphene Landau levels embedded in a deep sub-wavelength hyperbolic cavity, where light is confined into ultrasmall mode volumes. By studying the spectrum, we discuss the emergence of polaritons, and disentangle the contributions of resonant quantum vacuum effects from those of purely electrostatic interactions. Finally, we study the hybridization between magnetoplasmons and the cavity modes.
@@ I Introduction
Electromagnetic radiation is among the most versatile probes for investigating the properties of matter, and technological advancements enable the exploitation of various spectral regions to interrogate material characteristics. For instance, X-ray diffraction techniques elucidate the atomic arrangement within crystals Compton , Raman spectroscopy characterizes the vibrational modes of molecules Raman , time-resolved pump-probe spectroscopy reveals ultrafast electron dynamics, and both terahertz and angle-resolved photoemission spectroscopy (ARPES) provide insights into low-energy excitations and electronic band structures Damascelli . After the seminal work of Purcell Purcell1946 , it is now well established that light can be tailored by means of resonant structures - or cavities - to enhance its interaction with matter. Indeed, the emergence of cavity quantum electrodynamics (C-QED) Raimond01 ; Haroche13 , where mirrors confine the electromagnetic field, has enabled the measurement of quantum properties of a single atom coupled to a single photon.
@@
Light-matter hybridization manifests itself as an avoided crossing between energy levels, known as the vacuum Rabi splitting $\Omega$Sanchez-Mondragon1983 ; agarwal1984 ; agarwal1985 ; Brune1996 ; Raimond01 , which in solid-state systems can be probed in transmission and reflection experimentsKhitrova2006 ; Todorov_PhysRevLett_2010 ; Todorov_PhysRevLett_2009 ; Scalari_Science_2012 ; Bitton . This phenomenon can be understood within both classical and quantum settingsDDB_arxiv_2023 but, in the context of C-QED, the quantum nature of the electric field $\hat{E}$ unveils a general correspondence between the vacuum Rabi splitting and quantum vacuum fluctuations of the electric field, $\Omega\sim\sqrt{\langle\hat{E}^{2}\rangle}$ dalibard1982 ; ciuti2005 ; DDB_arxiv_2023 . A promising strategy to enhance vacuum fluctuations (and consequently the coupling) is to confine the electromagnetic field to a small volume $V$chang2006 ; Kockum19 ; tame2013 , as the amplitude of fluctuations scales as $1/\sqrt{V}$. While Fabry-Pérot cavities are often usedRaimond01 ; Haroche13 , the minimal achievable volume in these objects is limited by the so-called diffraction limit, $V\geq\lambda_{\rm d}^{3}$, where $\lambda_{\rm d}\equiv\lambda_{0}/2$ and $\lambda_{0}$ is the wavelength in free space. This limit imposes a significant obstacle to the compression and impact of quantum vacuum fluctuations.
@@
Conversely, sub-wavelength cavities, made from an optically active electromagnetic environment, differ from optical cavities as they can confine light in a volume $V\ll\lambda_{\rm d}^{3}$, well below the diffraction limit chang2006 ; koppens2011 ; tame2013 ; bozhevolnyi2006 ; oulton2009 ; baumberg2019 ; tame2013 . Paradigmatic examples of sub-wavelength cavities are plasmonic systems and metamaterials Maier2007 . Plasmonic nanocavities exploit the excitation of surface plasmons at metal-dielectric interfaces to confine electromagnetic fields to volumes significantly smaller than the diffraction limit Schuller2010 ; Ciraci2012 ; Chikkaraddy2016 . While these systems can achieve extremely high compression factorsAlcaraz2018 ; Barra-Burillo2021 , they suffer from a low quality factor ($Q\approx 10$) stemming from the highly dissipative nature of metallic systems. Among these, split-ring resonators operating in the terahertz regimeMaissen2014 have been extensively used in solid-state cavities experiments, due to their integrability with high-quality two-dimensional electron gases. More generally, polaritonic systems serve as promising platforms for solid-state C-QED, as they can host deep sub-wavelength optical modes and offer great tunability dai2014 . While this approach is experimentally promising, most current theories are developed for Fabry-Pérot cavities, where excitations are described by a transverse vector potential $\hat{\bm{A}}$. These theories are applied to sub-wavelength cavities by using a phenomenological substitution to enhance the Rabi frequency with a compression factor $A=V/\lambda_{\rm d}^{3}$. The compression factor quantifies how much a sub-wavelength cavity surpasses the diffraction limit Keller2017 ; arwas2023 , reaching values of $A\approx 10^{-10}$. Given a Rabi frequency $\Omega$ derived from optical cavity theory, which scales as $\sim 1/\sqrt{V}$, the same quantity for a sub-wavelength cavity is often obtainedgao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 ; Dag2024 by substituting $\Omega\to\Omega/\sqrt{A}$, thus enhancing light-matter interactions by a factor up to $10^{5}$. This phenomenological procedure underlies striking predictions in sub-wavelength cavities, such as cavity-mediated superconductivity schlawin2019 , breakdown of topological integer quantum Hall effect arwas2023 , and cavity control of topological properties wang2019 ; nguyen2023 .
However, it has been shown andolina2024 that, in the generalized Coulomb gauge, sub-wavelength cavities characterized by a strong compression $A\ll 1$, cannot be described by a transverse vector potential $\hat{\bm{A}}$. Instead, they are more accurately described by a scalar potential $\hat{\phi}$, reflecting their electrostatic characterandolina2024 ; pantazopoulos2024 ; Muniain2024 . This underscores the need for a comprehensive theory of sub-wavelength cavity QED based on the coupling with a scalar potential $\hat{\phi}$ solving the Poisson equation. Developing this theory—and applying it to the specific problem of cavity QED of Landau levels in graphene—is the main goal of the present article.
@@
The excitations of a sub-wavelength cavity are not transverse photons but longitudinal hybrid light-matter excitations such as plasmonsGPlasmonic and polaritonsPCavities . Recently, sub-wavelength cavities have been shown riolo2024 to alter the quasiparticle properties of Fermi liquids realized in weakly correlated 2D electron systems. Among sub-wavelength platforms, hyperbolic phonon-polariton (HPP) cavities dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore based on hexagonal boron nitride (hBN) offer an unparalleled balance of high compression factor ($A\approx 10^{-6}$) and impressive quality factors ($Q\approx 400$) HerzigSheinfux . Due to the strong anisotropy in its optical phonon spectrumcaldwell2014 , hBN exhibits an upper ($\sim 168-200$ meV) and a lower ($\sim 94-103$ meV) Reststrhalen band, where the in-plane and out-of-plane permittivity components are opposite in sign, leading to a hyperbolic polaritonic dispersion. The resulting ray-like nature of hBN mid-infrared phonon polaritons allows for extreme wavelength compression of light, enabling electromagnetic waves to be confined to deeply sub-wavelength scalesli2015 . On the other hand, the bulk nature of hBN phonon polaritons protects them from Ohmic loss mechanisms typical of metal-dielectric interfaces, enhancing the lifetime and making HPP cavities highly effective for mid-IR photonic applicationscaldwell2014 .
@@
In this paper, we present a fully non-local quantum theory to describe the C-QED of 2D electron systems coupled to sub-wavelength cavities, with the goal of deriving vacuum Rabi frequencies from a fully microscopic framework without relying on any phenomenological assumptions. To achieve this, we develop a quantization scheme for sub-wavelength cavities that enables the calculation of quantum vacuum fluctuations of the cavity electric field. This approach enables the use of standard methods and techniques characteristic of quantum optics and C-QED. While our theory is general for any sub-wavelength cavity, we apply it to an experimentally relevant scenario. Specifically, we investigate electrons in graphene subject to a strong perpendicular magnetic field ${\bm{B}}$ and coupled to the sub-wavelength modes of HPP cavity. This ${\bm{B}}$ field induces circular electron motion (cyclotron motion), leading to the emergence of quantized energy levels known as Landau levels (LLs)Giuliani_and_Vignale ; PrangeGirvin1987 . The excitations of HPP cavities made from hBN are typically around $0.1~{}{\rm eV}$ HerzigSheinfux , hence LL transitions can be tuned to this energy range by applying a magnetic field of a few Tesla. Therefore, if the coupling is higher than the losses, LLs transitions hybridize with resonant cavity modes, forming Landau-phonon-polaritons (LPhPs) hagenmuller2010 ; hagenmuller2012 ; Chirolli2012 ; Pellegrino2014 ; ashida2023 .
@@
To describe LPhPs, we develop a comprehensive microscopic model of quantum vacuum fluctuations in deep sub-wavelength HPP cavities. With our microscopic theory, we can predict the optimal cavity size to maximize light-matter coupling, measured by the vacuum Rabi frequencies, without resorting to any phenomenological compression factor. Importantly, the deep sub-wavelength regime allows cavity and matter excitations to hybridize at very high wavevectors, necessitating a fully non-local theoretical framework beyond the dipole approximation. Indeed, the maximal vacuum Rabi frequencies are obtained for finite momenta. Scattering-type near-field optical microscopy (s-SNOM) basov2016 , capable of probing collective modes at such high wavevectors, provides a promising platform to measure our predictions. The fully microscopic theory developed here can have profound implications for the control of materials in sub-wavelength cavities. As a major example, it will be interesting in the future to use our theory to study quantum transport in these 2D electron systems embedded in HPP cavities. Several experiments showed that sub-wavelength split-ring cavities affect the integer appugliese2022 ; enkner2024 and fractional enkner2024b quantum Hall, generating significant theoretical interest ciuti2021 ; arwas2023 ; bacciconi2023 ; rokaj2023 ; rokaj2022 . In principle, our framework can be applied to study how HPP cavities modify magnetotransport paravicini2019 and topology in general arwas2023 ; nguyen2023 ; wang2019 ; Zeno ; Dmytruk2022 ; Dmytruk2024 ; Dag2024 . In our case, as the collective magnetic excitations are strongly coupled to the cavity, a fully many-body treatment of quantum transport would be needed, rather than a single-particle approach. Since in sub-wavelength cavities the light-matter coupling occurs via the density operator, rather than via the current, these cavities circumvent the so-called “no-go theorems”andolina_prb_2019 ; andolina_epjp_2022 ; rzazewski_prl_1975 ; nataf_naturecommun_2010 and thus can, in principle, realize a superradiant phase transition. However, since the excitations of the cavity are encoded in the scalar potential $\hat{\phi}$, the nature of the phase transition would be ferroelectricdebernardis2018 ; rouse2023 , and would not lead to photon condensation.
@@
This paper is organized as follows. In Sect. II, we summarize the main concepts and key results of this work, which are presented in greater detail in the subsequent sections. In Section III, we develop a quantum theory of HPP cavities. Sect. IV provides an overview of the physics of LLs in graphene. In Sect. V, we delve into the details of light-matter interactions and analyze the resulting polaritonic spectrum. Additional technical details regarding the electrostatic properties of hexagonal boron nitride are provided in Appendix A, while the density-density response function of Landau levels in graphene is reviewed in Appendix B.
@@ II Summary of main results
In this article, we present our main findings as described in what follows. First, in Sect. II.1, we solve the Poisson equation with a frequency-dependent permittivity tensor within a sub-wavelength cavity to determine the scalar Green’s function. This solution enables us to quantize the scalar potential $\hat{\phi}$, expressing it in terms of the quantum modes of the cavity. This provides a theoretical framework for describing electromagnetic and quantum vacuum fluctuations within a sub-wavelength cavity. In Sect. II.2 we review the physics of LLs emerging in graphene subjected to a strong perpendicular magnetic field. Crucially, we highlight that, in sub-wavelength cavities, electrons couple through the density as they are coupled to a scalar potential $\hat{\phi}$. This in stark difference with traditional Fabry-Pérot cavities, where the light-matter interaction is implemented through the minimal coupling mechanism, which couples the electronic transverse current to the transverse vector potential. In Sect. II.3 we develop the C-QED of a 2D material embedded in a sub-wavelength cavity. By formulating and diagonalizing the Hopfield Hamiltonian, we study polaritonic states that arise from the hybridization of LL transitions with cavity modes. This also allows us to directly calculate, without invoking any phenomenological procedure, the vacuum Rabi frequencies of the system, showing the system achieves ultra-strong and super-strong coupling. While our theory is specifically applied to graphene embedded in an hBN cavity, the same theory can be readily applied to other 2D materials within sub-wavelength cavities. Finally, in Sect. II.4 we comment on the relation of theory with the previous literature.
@@ II.1 Theory of the Green’s function of HPP cavities
Generally, electromagnetic cavities are fully characterized by the solutions of Maxwell’s equationsJackson . In sub-wavelength cavities the most important contribution to the description of the electromagnetic properties is provided by the solution of the frequency-dependent Poisson equation in the presence of an external chargeandolina2024 ; riolo2024 :
@@
The cavity geometry that we consider is shown in Fig. 1. We further assume translational invariance along the $\hat{\bm{x}}$ and $\hat{\bm{y}}$ directions that denote 2D positions in this plane with the vector $\bm{r}_{\parallel}=x\hat{\bm{x}}+y\hat{\bm{y}}$. In the direction perpendicular to this plane, i.e. the $\hat{\bm{z}}$ direction and for $|z|\leq L_{z}/2$, we assume that space is filled by a hBN layer of thickness $L_{z}$ dai2014 ; caldwell2014 ; HerzigSheinfux ; li2015 ; Autore , characterized by a uniaxial permittivity tensortomadin2015 ; riolo2024 $\bm{\epsilon}(\bm{r},\omega)=\text{diag}(\epsilon_{x}(\bm{r},\omega),\epsilon_{x}(\bm{r},\omega),\epsilon_{z}(\bm{r},\omega))$. This dielectric tensor is initially assumed to be real, with its imaginary part introduced later through the substitution $\omega\to\omega+\mathrm{i}\eta$. While this approach is not generally required, it is particularly advantageous in this context, as it allows us to quantize the field first and incorporate losses subsequently. Further details on this tensor are reported in Appendix A. For $|z|>L_{z}/2$, we assume the presence of two perfect conductors, acting as mirrors. Consequently, the boundary conditions for the potential inside hBN ( $|z|\leq L_{z}/2$) are given by Jackson : $\phi(\bm{r}_{\parallel},z=\pm L_{z}/2)=0$.
@@
where $\bm{q}_{\parallel}$ is the wavevector in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane, ${q}_{\parallel}$ is its modulus, and $q_{z,n_{z}}\equiv(2n_{z}+1)\pi/L_{z}$. Eq. (LABEL:eqn:green_function_hBN_expansion_Main) represents a key theoretical result, as it shows that the scalar Green’s function can be expanded into distinct modes, which can then be quantized systematically. Notice that the Green’s function depends only on the modulus ${q}_{\parallel}$ as the system is invariant under rotations in the $\hat{\bm{x}}-\hat{\bm{y}}$ plane. Fig. 2(a) displays the spectrum $-{\rm Im}~{}g^{(0)}({q}_{\parallel},0,0,\omega)$ of the HPP cavity, which quantifies the absorption of the cavity at the position $z=z^{\prime}=0$, where we will place the 2D material. The scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ exhibits poles when the condition ${\epsilon}_{z}(\omega)q_{z,n_{z}}^{2}+{\epsilon}_{x}(\omega)q_{\parallel}^{2}=0$ is satisfied. This occurs only in the lower ($\omega_{z}^{\rm T}<\omega<\omega_{z}^{\rm L}$) and upper ($\omega_{x}^{\rm T}<\omega<\omega_{x}^{\rm L}$) Reststrahlen bands. The longitudinal and transverse phonon frequencies, $\omega_{i}^{\rm L}$ and $\omega_{i}^{\rm T}$, are marked by black dotted lines in Fig. 2(a). We denote the poles of the bare scalar Green’s function with the symbol $\omega_{{q}_{\parallel},n_{z},s}$, where $s=+1$ ($s=-1$) refers to the upper (lower) Reststrahlen band. There are important differences and similarities between the scalar Green’s function in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) for HPP cavities and the full dyadic Green’s function for Fabry-Pérot cavities andolina2024 ; Charlie_Thesis . Both functions share key features, such as poles corresponding to mode frequencies and a prefactor of $1/L_{z}$, which encodes the compression mechanism. However, in both cases, the dependence on $L_{z}$ is non-trivial, as the mode frequencies $\omega_{{q}_{\parallel},n_{z},s}$ themselves depend on the cavity length $L_{z}$. In Fabry-P’erot cavities, the poles, which are associated with the energies of cavity photons, appear only in the vectorial part of the dyadic Green’s function andolina2024 , while the scalar part only represents the Coulomb potential modified by the image charges induced in the mirrors. Conversely, in HPP cavities, poles corresponding to the phonon-polariton energies are present in both the scalar and vectorial components of the dyadic Green’s function. Nevertheless, due to the sub-wavelength nature of HPP cavity modes ($\lambda_{\rm p}=2\pi/q_{\parallel}\ll\lambda_{0}$) the scalar Green’s function $g^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ is the dominant contribution to the system’s electromagnetic properties andolina2024 .